In a p-n junction in Si, Nd = Na = 10^22 m^−3.
Calculate the built in potential for Si at
300 K if ni for Si at 300K=10^16m^−3.
1
Expert's answer
2020-04-17T10:19:12-0400

Use this


"\phi=\frac{kT}{q}ln(\frac{N_a\cdot N_d}{n_i^2})""\phi=0.0259\cdot ln(\frac{10^{22}\cdot 10^{22}}{10^{16}\cdot 10^{16}})=0.77 (V)"